OpenStudy (anonymous):

HELP NOW PLEASE!!! Find the X intercept. 5X - 6Y = 30

OpenStudy (anonymous):

what do x and y equal

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

oh ok um....

OpenStudy (the_fizicx99):

To find the x intercept, set y for 0 and solve. 5x - 6y = 30 5x - 6(0) = 30 5x = 30 /5 /5 x = 6 or (6,0)

OpenStudy (anonymous):

what are the answer options